OpenStudy (anonymous):

Find a cubic function f(x) = ax^3 + bx^2 + cx + d that has a local maximum value of 2 at x = −2 and a local minimum value of 0 at x = 1.

OpenStudy (cruffo):

Given: f(-2) = -8a + 4b - 2c + d = 2 f(1) = a + b + c + d = 0 Find the first derivative: f'(x) = 3ax^2 + 2bx + c Know critical points: f'(-2) = 12a - 4b + c = 0 f'(1) = 3a +2b + c = 0 So you have the system of equations: -8a + 4b - 2c + d = 2 a + b + c + d = 0 12a - 4b + c = 0 3a +2b + c = 0

OpenStudy (anonymous):

how do i find a, b, c, d?

OpenStudy (anonymous):